Use the table below for Problems 22 - 25. The debt and equity columns describe different possible capital structures A-G for Company XYZ. The Method 1-3 columns indicate the WACC for capital structures A-G and are consistent with the Modigliani and Miller Propositions. Structure Total Assets 8000 Equity Method 1 16.13% 13.84% 13.20% 12.24% 11.00% 11.20% 13.00% Method 2 5.49% 6.28% 9.47% 9.72% 9.60% 10.73% 13.00% Method 3 13.00% 13.00% 13.00% 1200 1000 800 600 400 400 400 400 400 400 400 13.00% 13.00% 22. The market risk premium is 10% and the risk free rate is 4.0%. Using the CAPM, calculate the (unlevered) beta for Company XYZ: B. 0.90 C. 1.00 D. 1.27 E. 1.30 23. Which method(s) imply that the tax shield from debt can at least sometimes reduce WACC? A. Method 1 B. Method 2 C. Method 3 D. Method 1 & Method 2 E. Method 1 & Method 3 24. Assume Company XYZs currently has $1000 in total assets and the higher the firms debt level, the higher the risk of bankruptcy. If the CFO is looking to maximize firm value, Company XYX should: A. borrow an additional $400 in debt B. repay or retire $400 in debt C borrow an additional $200 in debt D. repay or retire $200 in debt make no changes to the capital structure 25. Assume Company XYZs currently has capital Structure C under Method 1 where the WACC is 13.20% based on an effective tax rate of 30% and a cost of debt of 18%. There are 200 shares currently outstanding. The company just reported (T-0) FCFF of 300 and FCFE of 250. All cash flows are expected to grow at 20% for each of the next three years, then at 6% beginning in year four (T-4) in perpetuity (or indefinitely). If you were an activist investor trying to gain control of Company XYZ, what is the most you would pay per share for the hostile takeover? A. $6.97 B. $9.62 C. $11.55 D. $15.49 E $31.36

***Would especially like help with #25***
Answer #1

Since you have requested help especially or question no. 25, I am solving question no. 25.

Please see the table below:

T

0

1

2

3

FCFF (= FCFF0 x (1 + g))

300.00

360.00

432.00

518.40

Growth rate (g)

20%

20%

20%

Terminal growth rate (gT)

6%

WACC (R)

13.20%

Terminal value of Cash flows (= FCFF3 x (1 + gT) / (R - gT)

7,632.00

PV factor(=(1+R)(-T))

0.8834

0.7804

0.6894

PV of FCFF (= PV Factor x FCFF)

318.02

337.12

357.38

PV of Terminal Value (PV factor of year 3 x Terminal value)

5,261.37

Firm's Value (+Sum of PV of FCFF and PV of terminal value)

6,273.90

This is the maximum value of the firm. Nos. of shares outstanding, N = 200

Hence, maximum price per share = Firm's value / N = 6,273.90 / 200 = $ 31.37

Please choose option (E)